OpenStudy (anonymous):

2x + y = 7 5x + y = 9 Solve the system of equations. (-2/3, 25/3) (2/3, 17/3) 2/3, 25/3)

OpenStudy (jiteshmeghwal9):

2x+y=7 y=7-2x putting y into next equation 5x+(7-2x)=9 3x+7=9 3x=2 x=2/3

OpenStudy (jiteshmeghwal9):

y=7-2x put the value of x here & gt y :)

OpenStudy (anonymous):

?